### v1

@@ 4 Factorization of factors ${\mathbb{P}}(\omega^{\delta})$
If $\delta$ is an ordinal, $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa\geq\omega$, $\omega^{\delta}=\sum_{\xi<\kappa}\omega^{\delta_{\xi}}$, where $\delta_{\xi}\leq\delta_{\zeta}$, whenever $\xi<\zeta<\kappa$, and $\mathop{\rm cf}\nolimits(\omega^{\delta_{\xi}})\geq\kappa$, for all $\xi<\kappa$, then $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
Proof. Let ${\mathbb{L}}_{\xi}\cong\omega^{\delta_{\xi}}$, for $\xi<\kappa$, where $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Then $\omega^{\delta}\cong{\mathbb{L}}:=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$ and, by (d) and (f) of Fact 2.2 and Proposition 4.3(a), $\langle P(\omega^{\delta})\setminus{\mathcal{I}}_{\omega^{\delta}},\subset_{\omega^{\delta}}\rangle=\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}({\mathbb{L}})=\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$. So, we show that $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$. For $A\subset L$ let $\mathop{\rm supp}\nolimits(A)=\{\xi<\kappa:L_{\xi}\cap A\neq\emptyset\}$. By Proposition 4.6(c) we have
@@
and, since $S^{\zeta}_{A}\subset\mathop{\rm supp}\nolimits(A)$, for $\zeta<\kappa$,
@@
We show that $f:\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where $f(S)=\bigcup_{\xi\in S}L_{\xi}$, for $S\in[\kappa]^{\kappa}$. First, if $S\in[\kappa]^{\kappa}$ and $\zeta<\kappa$, then $|S\setminus\zeta|=\kappa$ and for $\xi\in S\setminus\zeta$ we have $\delta_{\zeta}\leq\delta_{\xi}$, which gives $\omega^{\delta_{\zeta}}\hookrightarrow\omega^{\delta_{\xi}}\cong L_{\xi}=L_{\xi}\cap f(S)$ and, hence, $\xi\in S^{\zeta}_{f(S)}$. Thus $[\kappa]^{\kappa}\ni S\setminus\zeta\subset S^{\zeta}_{f(S)}$, for each $\zeta<\kappa$, and by (21) we have $f(S)\in P(L)\setminus{\mathcal{I}}$. So, $f$ is a mapping from $[\kappa]^{\kappa}$ to $P(L)\setminus{\mathcal{I}}$ indeed. Let $S,T\in[\kappa]^{\kappa}$.
@@
(ce2) If $S\perp T$, then $|\mathop{\rm supp}\nolimits(f(S)\cap f(T))|=|\mathop{\rm supp}\nolimits(\bigcup_{\xi\in S\cap T}L_{\xi})|=|S\cap T|<\kappa$ and by (22) $f(S)\cap f(T)\in{\mathcal{I}}$, that is $f(S)\perp_{\subset_{\mathcal{I}}}f(T)$. If $S\not\perp T$, then $S\cap T\in[\kappa]^{\kappa}$ and $f(S)\cap f(T)=\bigcup_{\xi\in S\cap T}L_{\xi}=f(S\cap T)\in P(L)\setminus{\mathcal{I}}$ and, hence, $f(S)\not\perp_{\subset_{\mathcal{I}}}f(T)$.
@@
By recursion we construct a sequence $\langle\alpha_{\xi}:\xi<\kappa\rangle$ such that for each $\xi<\kappa$ we have
@@
So we can take $\alpha_{0}=\min S^{0}_{A}$ and $\alpha_{\zeta}=\min(S^{\zeta}_{A}\setminus\{\alpha_{\xi}:\xi<\zeta\})$, for $\zeta<\kappa$. Now we have $S:=\{\alpha_{\zeta}:\zeta<\kappa\}\in[\kappa]^{\kappa}$ and, by (25), $\{\alpha_{\xi}:\xi\geq\zeta\}\subset S^{\zeta}_{A}$, which implies that $S\subset^{*}S^{\zeta}_{A}$, for all $\zeta<\kappa$. By (23), $S$ is a reduction of $A$ to $\langle[\kappa]^{\kappa},\subset^{*}\rangle$. $\Box$
@@
If $\delta=\theta+\kappa$, where $\theta$ is an ordinal and $\theta\geq\kappa=\mathop{\rm cf}\nolimits(\delta)>\mathop{\rm cf}\nolimits(\theta)=:\lambda\geq\omega$, then $\langle[\lambda]^{\lambda},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
@@
Now $K:=\{\xi_{\nu}:\nu<\lambda\}\in[\kappa\setminus\eta]^{<\kappa}$ and $\sum_{\nu<\lambda}\omega^{\delta_{\nu}}\hookrightarrow\bigcup_{\nu<\lambda}L_{\xi_{\nu}}\cap A$, that is $\omega^{\theta}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$. So (26) is true. (b) follows from (a). $\Box$
Let $\{I_{\nu}:\nu<\lambda\}\subset[\kappa]^{\kappa}$ be a partition of $\kappa$. For $\nu<\lambda$ and $\xi\in I_{\nu}$, let $L^{\prime}_{\xi}\subset L_{\xi}$ be the initial part of ${\mathbb{L}}_{\xi}$ such that ${\mathbb{L}}_{\xi}^{\prime}\cong\omega^{\delta_{\nu}}$ and let $C=\bigcup_{\nu<\lambda}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}$. Then for each $\nu<\lambda$ the set $S^{\nu}_{C}=\{\xi<\kappa:\omega^{\delta_{\nu}}\hookrightarrow L_{\xi}\cap C\}=\bigcup_{\mu\geq\nu}I_{\mu}$ is of size $\kappa$ and, by Claim 4.9(a), $C\in{\mathbb{P}}(\omega^{\delta})$. Thus ${\mathbb{P}}(C)\cong{\mathbb{P}}(\omega^{\delta})$ and we prove that $f:\langle[\lambda]^{\lambda},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(C)$, where for $S\in[\lambda]^{\lambda}$
@@
If $\delta>0$ is an ordinal and $\mathop{\rm cf}\nolimits(\delta)=\kappa$, then regarding ${\mathbb{P}}(\omega^{\delta})$ we have
@@
    then $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed and $\langle[\omega]^{\omega},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$;
@@
    then $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed and $\langle[\omega]^{\omega},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$;
@@
    then $\langle[\lambda]^{\lambda},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$;
@@
    then $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$;
@@
    then $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
In cases (C), (D) and (E) forcing ${\mathbb{P}}(\omega^{\delta})$ collapses (at least) $\omega_{2}$ to $\omega$.
Proof. Case (A) (a) Let $\delta=\theta+1$; for $\theta=0$ we have $\omega^{\delta}=\omega$ and $\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega)=\langle[\omega]^{\omega},\subset^{*}\rangle$. If $\theta\geq 1$, then by Lemma 4.4(b) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}=\sum_{\omega}\omega^{\theta}$. Since $\theta\geq 1$ we have $\mathop{\rm cf}\nolimits(\omega^{\theta})\geq\omega$; so, by Theorem 4.7 we have $\langle[\omega]^{\omega},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
If $\mathop{\rm cf}\nolimits\delta=\omega$, let $\delta=\sup\{\delta_{i}:i<\omega\}$, where $\langle\delta_{i}:i<\omega\rangle$ is an increasing sequence of non-zero ordinals. Then by Lemma 4.4(a) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}=\sum_{i<\omega}\omega^{\delta_{i}}$. For $i<\omega$ we have $\delta_{i}>0$ and, hence, $\mathop{\rm cf}\nolimits(\omega^{\delta_{i}})\geq\omega$. Thus, by Theorem 4.7 we have $\langle[\omega]^{\omega},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
(b) By Lemma 4.4 we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}\cong\sum_{i<\omega}\omega^{\delta_{i}}=\sum_{i<\omega}L_{i}=:L$. Thus $\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}(L)=\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where :=${\mathcal{I}}:=\{A\subset L:L\not\hookrightarrow A\}$. For $\delta>1$ as in (a) we conclude that $\mathop{\rm cf}\nolimits(\omega^{\delta_{i}})\geq\omega$, for all $i<\omega$, and, by Proposition 4.6(c), for $A\subset L$ we have
@@
For $A_{n}\in P(L)\setminus{\mathcal{I}}$, $n\in\omega$, where $\dots\subset_{\mathcal{I}}A_{1}\subset_{\mathcal{I}}A_{0}$, we will construct $A\in P(L)\setminus{\mathcal{I}}$ such that $A\subset_{\mathcal{I}}A_{n}$, for all $n\in\omega$. Since ${\mathcal{I}}$ is an ideal we have $P(L)\setminus{\mathcal{I}}\ni B_{i}:=\bigcap_{n\leq i}A_{n}\subset A_{i}$, so we can assume that $A_{0}\supset A_{1}\supset A_{2}\supset\dots$. By recursion we define the sequences $\langle j_{i}:i\in\omega\rangle$ and $\langle f_{i}:i\in\omega\rangle$ such that for each $i\in\omega$ we have
@@
Case (B) (a) follows from Theorem 4.8, for $\lambda=\omega$.
(b) By Lemma 4.1, $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$. Clearly we have $\omega^{\delta}=\omega^{\theta}\omega^{\kappa}=\sup\{\omega^{\delta_{n}}:n\in\omega\}\kappa$. Let ${\mathbb{L}}=\langle L,<\rangle=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$, where ${\mathbb{L}}_{\xi}=\langle L_{\xi},<_{\xi}\rangle\cong\omega^{\theta}$, for $\xi<\kappa$, and $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Let =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$ and, for $A\subset L$ and $n\in\omega$, let
@@
Case (C) The statement follows from Theorem 4.8, for $\lambda>\omega$.
@@
Case (D) Clearly we have $\mathop{\rm cf}\nolimits(\delta)=\kappa$ so, by Lemma 4.4(a), $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$. In addition we have $\omega^{\delta}=\omega^{\theta+\kappa}=\omega^{\theta}\kappa=\sum_{\kappa}\omega^{\theta}$ and, by Lemma 4.4(a) again, $\mathop{\rm cf}\nolimits(\omega^{\theta})=\mathop{\rm cf}\nolimits(\theta)\geq\kappa$. Thus, by Theorem 4.7, $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.

### v2

@@ 3 Factorization of ${\mathbb{P}}(\alpha)$
If $\delta$ is an ordinal, $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa\geq\omega$, $\omega^{\delta}=\sum_{\xi<\kappa}\omega^{\delta_{\xi}}$, where $\delta_{\xi}\leq\delta_{\zeta}$, whenever $\xi<\zeta<\kappa$, and $\mathop{\rm cf}\nolimits(\omega^{\delta_{\xi}})\geq\kappa$, for all $\xi<\kappa$, then ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
Proof. Let ${\mathbb{L}}_{\xi}\cong\omega^{\delta_{\xi}}$, for $\xi<\kappa$, where $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Then $\omega^{\delta}\cong{\mathbb{L}}:=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$ and, by Fact 2.1(d) and Fact 4.3, $\langle P(\omega^{\delta})\setminus{\mathcal{I}}_{\omega^{\delta}},\subset_{\omega^{\delta}}\rangle=\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}({\mathbb{L}})=\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$. For $A\subset L$ by Proposition 4.6(c) we have
@@
and, since $S^{\zeta}_{A}\subset\mathop{\rm supp}\nolimits(A):=\{\xi<\kappa:L_{\xi}\cap A\neq\emptyset\}$, for $\zeta<\kappa$,
@@
By Fact 2.2 it is sufficient to show that $f:\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where $f(S)=\bigcup_{\xi\in S}L_{\xi}$, for $S\in[\kappa]^{\kappa}$. First, if $S\in[\kappa]^{\kappa}$ and $\zeta<\kappa$, then $|S\setminus\zeta|=\kappa$ and for $\xi\in S\setminus\zeta$ we have $\delta_{\zeta}\leq\delta_{\xi}$, which gives $\omega^{\delta_{\zeta}}\hookrightarrow\omega^{\delta_{\xi}}\cong L_{\xi}=L_{\xi}\cap f(S)$ and, hence, $\xi\in S^{\zeta}_{f(S)}$. Thus $[\kappa]^{\kappa}\ni S\setminus\zeta\subset S^{\zeta}_{f(S)}$, for each $\zeta<\kappa$, and by (12) we have $f(S)\in P(L)\setminus{\mathcal{I}}$. So, $f$ maps $[\kappa]^{\kappa}$ to $P(L)\setminus{\mathcal{I}}$. Let $S,T\in[\kappa]^{\kappa}$.
@@
(ce2) If $S\perp T$, then we have $|\mathop{\rm supp}\nolimits(f(S)\cap f(T))|=|\mathop{\rm supp}\nolimits(\bigcup_{\xi\in S\cap T}L_{\xi})|=|S\cap T|<\kappa$ and by (13) $f(S)\cap f(T)\in{\mathcal{I}}$, that is $f(S)\perp_{\subset_{\mathcal{I}}}f(T)$. If $S\not\perp T$, then $S\cap T\in[\kappa]^{\kappa}$ and $f(S)\cap f(T)=\bigcup_{\xi\in S\cap T}L_{\xi}=f(S\cap T)\in P(L)\setminus{\mathcal{I}}$ and, hence, $f(S)\not\perp_{\subset_{\mathcal{I}}}f(T)$.
@@
By recursion we construct a sequence $\langle\alpha_{\xi}:\xi<\kappa\rangle$ such that for each $\xi<\kappa$
@@
So, taking $\alpha_{0}=\min S^{0}_{A}$ and $\alpha_{\zeta}=\min(S^{\zeta}_{A}\setminus\{\alpha_{\xi}:\xi<\zeta\})$, for $\zeta<\kappa$, we have $S:=\{\alpha_{\zeta}:\zeta<\kappa\}\in[\kappa]^{\kappa}$ and, by (16), $\{\alpha_{\xi}:\xi\geq\zeta\}\subset S^{\zeta}_{A}$, which implies that $S\subset^{*}S^{\zeta}_{A}$, for all $\zeta<\kappa$. By (14), $S$ is a reduction of $A$ to $\langle[\kappa]^{\kappa},\subset^{*}\rangle$. $\Box$
@@
If $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa=\mathop{\rm cf}\nolimits(\delta)>\mathop{\rm cf}\nolimits(\theta)=:\lambda\geq\omega$, then ${\mathcal{P}}_{\lambda}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
@@
Now $K:=\{\xi_{\nu}:\nu<\lambda\}\in[\kappa\setminus\eta]^{<\kappa}$ and $\sum_{\nu<\lambda}\omega^{\delta_{\nu}}\hookrightarrow\bigcup_{\nu<\lambda}L_{\xi_{\nu}}\cap A$, that is $\omega^{\theta}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$. So (17) is true. Statement (b) follows from (a). $\Box$
Let $\{I_{\nu}:\nu<\lambda\}\subset[\kappa]^{\kappa}$ be a partition of $\kappa$. For $\nu<\lambda$ and $\xi\in I_{\nu}$, let $L^{\prime}_{\xi}\subset L_{\xi}$ be the initial part of ${\mathbb{L}}_{\xi}$ such that ${\mathbb{L}}_{\xi}^{\prime}\cong\omega^{\delta_{\nu}}$ and let $C=\bigcup_{\nu<\lambda}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}$. Then for each $\nu<\lambda$ the set $S^{\nu}_{C}=\{\xi<\kappa:\omega^{\delta_{\nu}}\hookrightarrow L_{\xi}\cap C\}=\bigcup_{\mu\geq\nu}I_{\mu}$ is of size $\kappa$ and, by Claim 4.9(a), $C\in{\mathbb{P}}(\omega^{\delta})$. Thus ${\mathbb{P}}(C)\cong{\mathbb{P}}(\omega^{\delta})$ and by Fact 2.2 it is sufficient to prove that $f:\langle[\lambda]^{\lambda},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(C)$, where for $S\in[\lambda]^{\lambda}$
@@
If $\delta>0$ is an ordinal and $\mathop{\rm cf}\nolimits(\delta)=\kappa$, then for ${\mathbb{P}}(\omega^{\delta})$ we have
@@
    then ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed;
@@
    then ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed;
@@
    then ${\mathcal{P}}_{\lambda}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$;
@@
    then ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$;
@@
    then ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
In cases (A) and (B) we have ${\mathbb{P}}(\omega^{\delta})\equiv_{forc}{\mathcal{P}}_{\omega}\ast\pi$, where $\pi$ is a ${\mathcal{P}}_{\omega}$-name for an $\omega$-distributive forcing. In other cases forcing ${\mathbb{P}}(\omega^{\delta})$ collapses (at least) $\omega_{2}$ to $\omega$.
Proof. (A) Let $\delta=\theta+1$. For $\theta=0$ we have $\omega^{\delta}=\omega$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega)={\mathcal{P}}_{\omega}$. If $\theta>0$, then by Lemma 4.4(b) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}=\sum_{\omega}\omega^{\theta}$. Since $\theta\geq 1$ we have $\mathop{\rm cf}\nolimits(\omega^{\theta})\geq\omega$; so, by Theorem 4.7, ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
If $\mathop{\rm cf}\nolimits\delta=\omega$, let $\delta=\sup\{\delta_{i}:i<\omega\}$, where $\langle\delta_{i}:i<\omega\rangle$ is an increasing sequence of non-zero ordinals. Then by Lemma 4.4(a) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}=\sum_{i<\omega}\omega^{\delta_{i}}$. For $i<\omega$ we have $\delta_{i}>0$ and, hence, $\mathop{\rm cf}\nolimits(\omega^{\delta_{i}})\geq\omega$. Thus, by Theorem 4.7, ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
Now we prove that the preorder $\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed. By Lemma 4.4 we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}\cong\sum_{i<\omega}\omega^{\delta_{i}}=\sum_{i<\omega}L_{i}=:L$. Thus $\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}(L)=\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where :=${\mathcal{I}}:=\{A\subset L:L\not\hookrightarrow A\}$. For $\delta>1$ as in (a) we conclude that $\mathop{\rm cf}\nolimits(\omega^{\delta_{i}})\geq\omega$, for all $i<\omega$, and, by Proposition 4.6(c), for $A\subset L$ we have
@@
For $A_{n}\in P(L)\setminus{\mathcal{I}}$, $n\in\omega$, where $\dots\subset_{\mathcal{I}}A_{1}\subset_{\mathcal{I}}A_{0}$, we construct $A\in P(L)\setminus{\mathcal{I}}$ such that $A\subset_{\mathcal{I}}A_{n}$, for all $n\in\omega$. Since ${\mathcal{I}}$ is an ideal we have $P(L)\setminus{\mathcal{I}}\ni B_{i}:=\bigcap_{n\leq i}A_{n}\subset A_{i}$, so we can assume that $A_{0}\supset A_{1}\supset A_{2}\supset\dots$. By recursion we define the sequences $\langle j_{i}:i\in\omega\rangle$ and $\langle f_{i}:i\in\omega\rangle$ such that for each $i\in\omega$ we have
@@
(B) By Theorem 4.8 (for $\lambda=\omega$) we have ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
By Lemma 4.1, $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$. Clearly we have $\omega^{\delta}=\omega^{\theta}\omega^{\kappa}=\sup\{\omega^{\delta_{n}}:n\in\omega\}\kappa$. Let ${\mathbb{L}}=\langle L,<\rangle=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$, where ${\mathbb{L}}_{\xi}=\langle L_{\xi},<_{\xi}\rangle\cong\omega^{\theta}$, for $\xi<\kappa$, and $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Let =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$ and, for $A\subset L$ and $n\in\omega$, let
@@
(C) The statement follows from Theorem 4.8 (for $\lambda>\omega$).
@@
(D) Clearly we have $\mathop{\rm cf}\nolimits(\delta)=\kappa$ so, by Lemma 4.4(a), $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$. In addition we have $\omega^{\delta}=\omega^{\theta+\kappa}=\omega^{\theta}\kappa=\sum_{\kappa}\omega^{\theta}$ and, by Lemma 4.4(a) again, $\mathop{\rm cf}\nolimits(\omega^{\theta})=\mathop{\rm cf}\nolimits(\theta)\geq\kappa$. Thus, by Theorem 4.7, ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.